### 05 Mar 2026 v1

@@ Abstract
We present guidelines for deriving new Nitsche Finite Element Methods to enforce equality and inequality constraints that act on the value of the unknown mechanical quality. We first formulate the problem as a stabilized finite element method for the saddle point formulation where a Lagrange multiplier enforces the underlying constraint. The Nitsche method is then presented in a general minimization form, suitable for nonlinear finite element methods and allowing straightforward computational implementation with automatic differentation. To validate these ideas, we present Nitsche formulations for a range of problems in solid mechanics and give numerical evidence of the convergence rates of the Nitsche method.
@@ 1 Introduction
In the 1970s, Nitsche [nitsche] introduced a method for enforcing the Dirichlet boundary conditions in the weak formulation of the finite element method. The benefits of the Nitsche method over its alternatives, the penalty method [courant1943variational] and the mixed Lagrange multiplier method [babuvska1973finite], include consistency, symmetry and positive definiteness, stability, and conditioning; cf. [gustafsson2025]. Hence, the ideas of Nitsche have later found applications also in domain decomposition [becker2003finite], discontinuous Galerkin methods [hansbo2002discontinuous], unfitted finite element and fictitious domain methods [burman2015cutfem], Robin boundary conditions [juntunen2009nitsche], Kirchhoff plates [gustafsson2021nitsche], elastic contact problems [chouly2013nitsche], friction problems [chouly2014adaptation], obstacle problems [gustafsson2017mixed], elastoplastic torsion problems [chouly2023nitsche], and plate contact problems [gustafsson2017nitsche], to name a few.
In modern literature, the Nitsche method is often presented as a consistency correction to the penalty method, i.e., to make sure that the exact solution satisfies the discrete weak form, by including additional terms; cf., e.g.,  [benzaken2024constructing, gustafsson2025]. This fails to answer to the obvious follow-up question: how to generalize the method to other problems and constraints? Consequently, we take a different viewpoint based on the saddle point formulation and present guidelines to derive optimal Nitsche methods for both equality and inequality constraints. These guiding principles apply to arbitrary constrained minimization problems, where the constraint acts on the values of the unknown. The resulting energy minimization form is particularly well-suited for solving nonlinear problems numerically. We follow the guidelines in practice by deriving and implementing novel methods for problems such as inequality boundary condition for plates, membrane-solid contact problem, and plate-plate contact problem.
@@
In this paper, built on these observations, we explain how to define and use general Nitsche methods, present examples of their adaptation to novel physical problems, and show numerical evidence of their convergence. A more careful numerical analysis, i.e., proofs of stability and error estimates, is beyond the scope of this work. The remainder of the paper is organized as follows. In Section 2, we review the original Nitsche method for the Dirichlet boundary condition $u=g$. In Section 3, we demonstrate how the method is modified to satisfy the inequality constraint $u\geq g$. In Section 4, we describe a general form of the Nitsche method and how it can be used to derive new methods. In Section 5, we present some methods that exist in the literature and are of this general form. In Section 6, we derive novel Nitsche methods from the general form and show numerical evidence of their performance.
@@ 2 Background: Dirichlet boundary conditions
This form can be compared to the penalty method which simply minimizes the penalized functional:
$J(u_{h})+\frac{1}{2\varepsilon}\int_{\partial\Omega}(u_{h}-g)^{2}.$
The penalty method approximates the Robin boundary condition
@@
which satisfies $u-g\rightarrow 0$ when $\varepsilon\rightarrow 0$. Unfortunately, small enough values of $\varepsilon$ lead to ill-conditioned linear systems [gustafsson2025].
@@ 3 Inequality constraints: Signorini’s problem
is often referred to as the scalar Signorini’s or Poisson–Signorini problem [gustafsson2019nitsche]. The corresponding continuous saddle point formulation of (16) is: find $(u,\lambda)\in H^{1}(\Omega)\times H^{-1/2}(\partial\Omega)$, $\lambda\geq 0$, such that
@@
for every $(v_{h},\mu_{h})\in V_{h}\times Q_{h}$, $\mu_{h}\geq 0$. The second inequality implies that $\lambda_{h}$ is the orthogonal $L^{2}$ projection of $\kappa\frac{\partial u_{h}}{\partial n}-\frac{1}{\gamma}(u_{h}-g)$ onto the set of nonnegative functions in $Q_{h}$ [brezis2011functional, Theorem 5.2]. Since $Q_{h}$ is arbitrary, we may choose $Q_{h}=L^{2}(\partial\Omega)$ and then the orthogonal projection is given explicitly by the maximum operator, i.e.,
@@ 4 Structure of the method
We now generalize the method to arbitrary inequality constraints of the form $\beta(u_{h})\geq 0$. The method (20) has the following main ingredients: the energy of the source problem, i.e., $J$; the constraint to impose, i.e., $\beta(u_{h})=u_{h}-g\geq 0$; the corresponding definition of the continuous Lagrange multiplier, i.e., $\lambda(u_{h})=\kappa\frac{\partial u_{h}}{\partial n}$; and the correct scaling of the stabilization parameter $\gamma$ with respect to the mesh parameter $h$ and any material parameters, i.e., $\gamma(h,\kappa)|_{K}=\tfrac{\alpha h_{K}}{\kappa}$. In conclusion, the general form of the Nitsche method can be written as follows: find $u_{h}\in V_{h}$ that minimizes the functional
@@
The final method is formulated as a minimization problem instead of a variational one as it is the starting point for our implementation of the method; see Section 6 and Appendix A. In particular, the calculation of the second functional derivative of (21) is performed automatically in our computational implementation of the Nitsche method.
@@
    A constraint $\beta=\beta(u_{h})\geq 0$ and the subset $\Gamma\subset\Omega$ where the constraint can become active.
@@ 6 Novel methods
For each problem, we present the Nitsche formulation and numerically investigate its convergence rate. The numerical results are computed using scikit-fem [gustafsson2020scikit] which is used to solve the nonlinear variational problems using Newton’s method. For completeness, we describe how to manipulate the energy minimization problems for Newton’s method in Appendix A. The programming implementation utilizes automatic differentiation with JAX [bradbury2021jax] so that it is not necessary to manually calculate the functional derivatives involved in Newton’s method. The source code is available in [sourcepackage] for reproducing the numerical results.
@@ 6.1 Two-membrane contact problem
The problem was solved numerically over $\Omega=(0,1)^{2}$ with the parameterization $g=0.05$, $f_{1}=1,f_{2}=0,\kappa_{1}=\kappa_{2}=1$ and $\alpha={1}\mathrm{e}{-2}$ using linear triangular elements. Note that the element-wise Laplacian for linear elements is equal to zero which simplifies the formulation. The numerical solution is displayed in Figure 1 and the approximated linear convergence rate in Figure 2.
Figure 2: Two membrane contact problem convergence rate follows the theoretical linear convergence in the $H^{1}$ norm for linear finite element basis.
As described earlier, we can revert to the penalty method by removing the terms depending on $\lambda((u_{1},u_{2}))$. However, the penalty method retains the optimal convergence rate only with linear elements, and to produce the same level of accuracy as Nitsche for quadratic or higher order elements we need to increase the power of $h$ in $\gamma$ [gustafsson2017finite]. This leads to higher condition numbers in the Newton iteration, see Figure 3 for an example in the quadratic case.
Figure 3: Condition numbers of the Jacobian in the Newton iterations (45) for the penalty and Nitsche variants of the two membrane contact problem with quadratic elements. The blue lines depict condition numbers of the penalty method and red lines of the Nitsche method, respectively. The markers distinguish problems of different sizes. Compared to the Nitsche method’s $\gamma=\alpha h^{2}$, the penalty method requires $\gamma=\alpha h^{3}$ to retain optimal convergence [gustafsson2017finite] and has thus consistently larger conditioning numbers with slower convergence.
@@ 6.2 Membrane against elastic solid
The problem was solved numerically using a gap of $g=0.1$, $\mu_{L}=\lambda_{L}=1$, $f=2$ and $\alpha={1}\mathrm{e}{-2}$ using linear hexahedral elements for the solid and linear quadrilateral elements for the membrane. The numerical solution is displayed in Figure 4 and the approximated linear convergence rate in Figure 5. The first principal stress for the solid is presented in Figure 6.
@@ 6.3 Plate against plate
The plate against plate problem follows very closely the two membrance contact problem with the difference of higher order derivatives in the strain energy. The energy of the source problem is
@@
where $\Delta_{h}^{2}$ denotes the element-wise biharmonic operator.
The problem was solved numerically over $\Omega=(0,1)^{2}$, a gap of $g=0.05$, $f_{1}=100$, $f_{2}=0$ and $\alpha={1}\mathrm{e}{-2}$ using triangular Morley elements [beirao2010posteriori]. Notice that the Morley element is quadratic and hence the element-wise biharmonic operator evaluates to zero. The numerical solution is displayed in Figure 7 and the approximated linear convergence rate in Figure 8.
@@ 6.4 Kirchhoff plate with inequality boundary condition
Let $\Omega=(0,1)^{2}$. For the plate corners problem [nazarov2012hinged], the energy with normalized units reads
@@
We consider the problem where $\Gamma=\partial\Omega$ and $\beta(u)=u\geq 0$ on $\Gamma$. Then, the scaling parameter is $\gamma(h)=\alpha h^{3}$ and the Lagrange multiplier is $\lambda(u)=K(u):=Q_{n}+\frac{\partial M_{ns}}{\partial s}$, where M=HuM=-Hu, $Q=\text{div}\>M$, $n$ is the normal vector and $s$ the tangent vector on $\Gamma$ and $K(u)$ is called the Kirchhoff shear force [gustafsson2021nitsche].
@@
$J(u_{h})+\int_{\Gamma}\frac{\alpha h^{3}}{2}\Big(K_{h}(u)-\frac{1}{\alpha h^{3}}u_{h}\Big)_{+}^{2}-\int_{\Gamma}\frac{\alpha h^{3}}{2}K_{h}(u)^{2},$ (39)
where $K_{h}$ is the element-wise evaluation of $K$.
@@ Acknowledgements
We wish to thank Rolf Stenberg for valuable comments on the manuscript.
@@ Appendix A Newton’s method
$\min_{u\in U}J(u)$ (40)
where $J:U\to\mathbb{R}$ represents the energy of the system. Let $v\in V$ denote an arbitrary perturbation to $u$. A solution $u\in U$ to (40) satisfies
@@
where $J^{\prime}:U\times V\to\mathbb{R}$ denotes the directional (Gateaux) derivative defined by
$J^{\prime}(u;v)=\frac{\partial J(u+\epsilon v)}{\partial\epsilon}\bigg|_{\epsilon=0}.$
The directional derivative is linear in $v$, but, in general, nonlinear in $u$.
Let $f=J^{\prime}$. We assume that the problem
$f(u;v)=0\quad\forall v\in V$
is nonlinear in $u$ and, hence, the standard Ritz–Galerkin method cannot be directly applied to obtain a linear system of equations. Consequently, we approach the problem as a root-finding problem and linearize $f$ to use the Newton’s method. More precisely, let $u_{0}\in U$ be an initial value such that the solution is written as $u=u_{0}+w$ for some $w\in V$. The perturbation $w$ is unknown, but can be approximated by discarding the higher order terms of the Taylor series of $f$:
$0=f(u;v)=f(u_{0}+w;v)\approx f(u_{0};v)+f^{\prime}(u_{0};v,w),$
where $f^{\prime}:U\times V\times V\to\mathbb{R}$ is the directional derivative of $f$.
A single step of the Newton’s method then reads: find $w\in V$ such that
$f^{\prime}(u_{0};v,w)=-f(u_{0};v)\quad\forall v\in V.$ (42)
Note that if $u_{0}\in U$ is fixed, then the problem is linear with respect to $w\in V$ and the standard Ritz–Galerkin method can be applied iteratively.
For the numerical scheme, we use the finite element method. Let $\mathcal{T}_{h}$ be a mesh of elements with maximum edge length $h>0$. The finite dimensional subspaces $U_{h}\subset U$ and $V_{h}\subset V$ contain element-wise polynomial functions of a given degree. The discrete problem then reads: given $u_{0,h}\in U_{h}$, find $w_{h}\in V_{h}$ such that
$f^{\prime}(u_{0,h};v_{h},w_{h})=-f(u_{0,h};v_{h})\quad\forall v_{h}\in V_{h}.$ (43)
The directional derivative of $f$ is computable as $u_{0,h}$ is given.
The equation (43) can be posed as a linear system of equations. Let $\{\phi_{i}\}_{i=1}^{n}$, $n=\dim(V_{h})$, be a finite element basis that span $V_{h}$. Then
$w_{h}=\sum_{j=1}^{n}\beta_{j}\phi_{j},$ (44)
where $\beta_{j}\in\mathbb{R}$ are the unknown degrees-of-freedom. Choosing $v_{h}=\phi_{i}$, we have
$\sum_{j=1}^{n}\beta_{j}f^{\prime}(u_{0,h};\phi_{i},\phi_{j})=-f(u_{0,h};\phi_{i}),$
and collecting all $i=1,\dots,n$ results in a linear system
$\bm{J}\bm{\beta}=\bm{f}$ (45)
where
$\bm{J}_{ij}=f^{\prime}(u_{0,h};\phi_{i},\phi_{j}),\quad\bm{f}_{i}=-f(u_{0,h};\phi_{i}),$
or, if the problem is given directly in the minimization form,
$\bm{J}_{ij}=J^{\prime\prime}(u_{0,h};\phi_{i},\phi_{j}),\quad\bm{f}_{i}=-J^{\prime}(u_{0,h};\phi_{i}).$
The direction (44) is then found from the solution $\bm{\beta}$.
To conclude, starting from an initial value $u_{0,h}\in U_{h}$, we find an approximation $u_{h}\in U_{h}$ such that $J^{\prime}(u_{h};v_{h})=0$ for all $v_{h}\in V_{h}$ through the iterative scheme $u_{k+1,h}=u_{k,h}+\alpha_{k}w_{k,h}$, where $0<\alpha_{k}\leq 1$ is a step size parameter. Under standard assumptions the method converges quadratically.

### 13 Apr 2026 v2

@@ Abstract
We present guidelines for deriving new Nitsche Finite Element Methods to enforce equality and inequality constraints that act on the value of the unknown mechanical quantity. We first formulate the problem as a stabilized finite element method for the saddle point formulation where a Lagrange multiplier enforces the underlying constraint. The Nitsche method is then presented in a general minimization form, suitable for adding constraints to nonlinear finite element methods and allowing straightforward computational implementation with automatic differentation. This extends the method beyond classical boundary condition enforcement. To validate these ideas, we present Nitsche formulations for a range of problems in solid mechanics and give numerical evidence of the convergence rates of the Nitsche method.
@@ 1 Introduction
In the 1970s, Nitsche [nitsche] introduced a method for enforcing the Dirichlet boundary conditions in the weak formulation of the finite element method. The benefits of the Nitsche method over its alternatives, the penalty method [courant1943variational] and the mixed Lagrange multiplier method [babuvska1973finite], include consistency, symmetry and positive definiteness, stability, and conditioning; cf. [gustafsson2025]. Hence, the ideas of Nitsche have later found applications also in domain decomposition [becker2003finite], discontinuous Galerkin methods [hansbo2002discontinuous], unfitted finite element and fictitious domain methods [burman2015cutfem], Robin boundary conditions [juntunen2009nitsche], Kirchhoff plates [gustafsson2021nitsche], elastic contact problems [chouly2013nitsche], friction problems [chouly2014adaptation], obstacle problems [gustafsson2017mixed], elastoplastic torsion problems [chouly2023nitsche], and plate contact problems [gustafsson2017nitsche, fabre2021nitsche], to name a few.
In modern literature, the Nitsche method is often presented as a consistency correction to the penalty method, i.e., to make sure that the exact solution satisfies the discrete weak form, by including additional terms; cf., e.g.,  [benzaken2024constructing, gustafsson2025]. This fails to answer to the obvious follow-up question: how to generalize the method to other problems and constraints? Consequently, we take a different viewpoint based on the saddle point formulation and present guidelines to derive optimal Nitsche methods for both equality and inequality constraints. These guiding principles apply to arbitrary constrained minimization problems, where an affine constraint acts on the values of the unknown. This extends the method beyond conventional boundary condition enforcement. The resulting energy minimization form is particularly well-suited for adding constraints to nonlinear problems. We follow the guidelines in practice by deriving and implementing novel methods for problems such as inequality boundary condition for plates, membrane-solid contact problem, and plate-plate contact problem.
@@
In this paper, built on these observations, we explain how to define and use general Nitsche methods, present examples of their adaptation to novel physical problems, and show numerical evidence of their convergence. A more careful numerical analysis, i.e., proofs of stability and error estimates for the general formulation, is beyond the scope of this work. The remainder of the paper is organized as follows. In Section 2, we review the original Nitsche method for the Dirichlet boundary condition $u=g$. In Section 3, we demonstrate how the method is modified to satisfy the inequality constraint $u\geq g$. In Section 4, we describe a general form of the Nitsche method and how it can be used to derive new methods. In Section 5, we present some methods that exist in the literature and are of this general form. In Section 6, we derive novel Nitsche methods from the general form and show numerical evidence of their performance.
@@ 2 Background: Dirichlet boundary conditions
We can formally compare the above to the penalty method (e.g., [gustafsson2025]) which is obtained by minimizing the penalized functional:
$J(u_{h})+\frac{1}{2\varepsilon}\int_{\partial\Omega}(u_{h}-g)^{2},\quad\varepsilon>0.$
The penalty method is known to approximate the Robin boundary condition
@@
which satisfies $u-g\rightarrow 0$ when $\varepsilon\rightarrow 0$.
@@ 3 Inequality constraints: Signorini’s problem
is often referred to as the scalar Signorini’s or Poisson–Signorini problem [gustafsson2019nitsche]. We note that problem (16) is semi-coercive: the bilinear form associated to the left-hand side controls only the $H^{1}$-seminorm, and the inequality constraint is not sufficient to restore full coercivity. We assume that $f$ and $g$ are such that a unique solution exists.
The corresponding continuous saddle point formulation of (16) is: find $(u,\lambda)\in H^{1}(\Omega)\times H^{-1/2}(\partial\Omega)$, $\lambda\geq 0$, such that
@@
for every $(v_{h},\mu_{h})\in V_{h}\times Q_{h}$, $\mu_{h}\geq 0$. The second inequality implies that $\lambda_{h}$ is the orthogonal $L^{2}$ projection of $\kappa\frac{\partial u_{h}}{\partial n}-\frac{1}{\gamma}(u_{h}-g)$ onto the set of nonnegative functions in $Q_{h}$ [brezis2011functional, Theorem 5.2]. Since $Q_{h}$ is arbitrary, we may choose $Q_{h}\subset L^{2}(\partial\Omega)$ such that the orthogonal projection is given explicitly by the maximum operator, i.e.,
@@ 4 Structure of the method
We now generalize the method to arbitrary inequality constraints of the form $\beta(u_{h})\geq 0$ where $\beta$ is an affine function. The method (20) has the following main ingredients: the energy of the source problem, i.e., $J$; the constraint to impose, i.e., $\beta(u_{h})=u_{h}-g\geq 0$; the corresponding definition of the continuous Lagrange multiplier, i.e., $\lambda(u_{h})=\kappa\frac{\partial u_{h}}{\partial n}$; and the correct scaling of the stabilization parameter $\gamma$ with respect to the mesh parameter $h$ and any material parameters, i.e., $\gamma(h,\kappa)|_{K}=\tfrac{\alpha h_{K}}{\kappa}$. In conclusion, the general form of the Nitsche method can be written as follows: find $u_{h}\in V_{h}$ that minimizes the functional
@@
The final method is formulated as a minimization problem instead of a variational one as it is the starting point for our implementation of the method; see Section 6 and A. In particular, the calculation of the second functional derivative of (21) is performed automatically in our computational implementation of the Nitsche method.
@@
    A constraint $\beta=\beta(u_{h})\geq 0$ and the subset $\Gamma\subset\Omega$ (or $\Gamma\subset\partial\Omega$) where the constraint can become active.
@@ 6 Novel methods
For each problem, we present the Nitsche formulation and numerically investigate its convergence rate. The numerical results are computed using scikit-fem [gustafsson2020scikit] which is used to solve the nonlinear variational problems using Newton’s method. For completeness, we describe how to manipulate the energy minimization problems for Newton’s method in A. The programming implementation utilizes automatic differentiation with JAX [bradbury2021jax] so that it is not necessary to manually calculate the functional derivatives involved in Newton’s method. The source code is available in [sourcepackage] for reproducing the numerical results.
@@ 6.1 Two-membrane contact problem
The problem was solved numerically over $\Omega=(0,1)^{2}$ with the parameterization $g=0.05$, $f_{1}=1,f_{2}=0,\kappa_{1}=\kappa_{2}=1$ and $\alpha=10^{-2}$ using both linear and quadratic triangular elements. Note that the element-wise Laplacian is equal to zero for linear elements and constant for quadratic elements, which simplifies the formulation. The linear element numerical solution is displayed in Figure 1 and the approximated convergence rates in Figure 2.
Figure 2: Two membrane contact problem convergence rates follow the theoretical linear in the $H^{1}$ norm for linear element basis. For quadratic basis, the convergence is slightly slower than quadratic. The latter convergence rate is limited by $C^{1.1}$ regularity, although the matching meshes might accelerate the empirical rate.
As described earlier, we can revert to the penalty method by removing the terms depending on $\lambda((u_{1},u_{2}))$. However, the penalty method retains the optimal convergence rate only with linear elements, and to produce the same level of accuracy as Nitsche for quadratic or higher order elements we need to increase the power of $h$ in $1/\epsilon$ [gustafsson2017finite]. This leads to higher condition numbers in the Newton iteration, see Figure 3 for an example in the quadratic case.
Figure 3: Condition numbers of the Jacobian in the Newton iterations (43) for the penalty and Nitsche variants of the two membrane contact problem with quadratic elements. The blue lines depict condition numbers of the penalty method and red lines of the Nitsche method, respectively. The markers distinguish problems of different sizes. Compared to the Nitsche method’s $\gamma=\alpha h^{2}$, the penalty method requires $1/\epsilon=\alpha h^{3}$ to retain optimal convergence [gustafsson2017finite] and has thus consistently larger conditioning numbers with slower convergence.
@@ 6.2 Membrane against elastic solid
The problem was solved numerically using a gap of $g=0.1$, $\mu_{L}=\lambda_{L}=1$, $f=2$ and $\alpha=10^{-2}$ using linear hexahedral elements for the solid and linear quadrilateral elements for the membrane. The numerical solution is displayed in Figure 4 and the approximated linear convergence rate in Figure 5. The first principal stress for the solid is presented in Figure 6.
@@ 6.3 Plate against plate
The plate against plate problem is a Kirchhoff–Love plate model with clamped plates, which follows very closely the two-membrane contact problem with the difference of higher order derivatives in the strain energy. The energy of the source problem is
@@
where $\Delta_{h}^{2}$ denotes the element-wise biharmonic operator. Note that the Nitsche method for plate obstacle problems were considered in [fabre2021nitsche], which suggests that the Kirchhoff model can be suitable as long as the transverse compression of the plate can be neglected.
The problem was solved numerically over $\Omega=(0,1)^{2}$, a gap of $g=0.05$, $f_{1}=100$, $f_{2}=0$ and $\alpha=10^{-2}$ using triangular Morley elements [beirao2010posteriori]. Notice that the Morley element is quadratic and hence the element-wise biharmonic operator evaluates to zero. The numerical solution is displayed in Figure 7 and the approximated linear convergence rate in Figure 8.
@@ 6.4 Kirchhoff plate with inequality boundary condition
Let $\Omega=(0,1)^{2}$. For the Kirchhoff plate with a simply-supported inequality boundary condition [nazarov2012hinged], the energy with normalized units reads
@@
We consider the problem where $\Gamma=\partial\Omega$ and $\beta(u)=u\geq 0$ on $\Gamma$. Then, the Lagrange multiplier is $\lambda(u)=V_{n}(u):=Q_{n}+\frac{\partial M_{ns}}{\partial s}$, where $V_{n}(u)$ is called the Kirchhoff shear force and
$M(u)=\frac{Ed^{3}}{12(1+\nu)}\left(K(u)+\frac{\nu}{1-\nu}(\text{tr}\,K(u))I\right),$
where $K(u)=-\nabla\nabla u$ is the curvature, $d$ denotes the plate thickness, $I$ is the identity tensor, and $E$ and $\nu$ are Young’s modulus and Poisson ratio, respectively. Further, $Q=\text{div}\>M$, $n$ is the normal vector and $s$ the tangent vector on $\Gamma$. For more details, we refer to [gustafsson2021nitsche]. It follows that the scaling parameter is $\gamma(h)=\alpha h^{3}$. As mentioned, we assume normalized units so that $\nu=0$ and $\frac{Ed^{3}}{12}=1$.
@@
$J(u_{h})+\int_{\Gamma}\frac{\alpha h^{3}}{2}\Big(V_{n,h}(u)-\frac{1}{\alpha h^{3}}u_{h}\Big)_{+}^{2}-\int_{\Gamma}\frac{\alpha h^{3}}{2}V_{n,h}(u)^{2},$ (39)
where $V_{n,h}$ is the element-wise evaluation of $V_{n}$.
@@ Acknowledgements
We wish to thank Rolf Stenberg and the anonymous referees for valuable comments on the manuscript.
@@ Appendix A Newton’s method
$\min_{u\in U}J(u),$ (40)
where $J:U\to\mathbb{R}$ represents the energy of the system. A solution $u\in U$ to (40) satisfies
@@
where $J^{\prime}(u;v)=\frac{\partial}{\partial\epsilon}J(u+\epsilon v)\big|_{\epsilon=0}$ denotes the directional (Gateaux) derivative. In general, (41) is nonlinear in $u$.
We linearize using Newton’s method. Let $u_{0}\in U$ be an initial value and write the solution as $u=u_{0}+w$ for some $w\in V$. Discarding the higher order terms of $J^{\prime}(u_{0}+w;v)$ gives the Newton step: find $w\in V$ such that
$J^{\prime\prime}(u_{0};v,w)=-J^{\prime}(u_{0};v)\quad\forall v\in V,$ (42)
where $J^{\prime\prime}:U\times V\times V\to\mathbb{R}$ is the second directional derivative. Since $u_{0}$ is fixed, the problem (42) is linear in $w$ and the Ritz–Galerkin method applies.
For the numerical scheme, let $\mathcal{T}_{h}$ be a mesh of elements with maximum edge length $h>0$ and let $U_{h}\subset U$, $V_{h}\subset V$ be finite element subspaces of element-wise polynomials of a given degree. Choosing a basis $\{\phi_{i}\}_{i=1}^{n}$ of $V_{h}$, the discrete Newton step (42) leads to the linear system
$\bm{J}\bm{\beta}=\bm{f},\quad\bm{J}_{ij}=J^{\prime\prime}(u_{0,h};\phi_{i},\phi_{j}),\quad\bm{f}_{i}=-J^{\prime}(u_{0,h};\phi_{i}),$ (43)
where $w_{h}=\sum_{j=1}^{n}\beta_{j}\phi_{j}$. Starting from $u_{0,h}\in U_{h}$, the iterative scheme $u_{k+1,h}=u_{k,h}+\alpha_{k}w_{k,h}$, with step size $0<\alpha_{k}\leq 1$, yields an approximation $u_{h}\in U_{h}$ satisfying $J^{\prime}(u_{h};v_{h})=0$ for all $v_{h}\in V_{h}$. Under standard assumptions the method converges quadratically.